A 4.0 kg block rests at the edge of a platform that is 20 m above level ground.  The block is launched horizontally with an initial velocity of 15 m/s. The time it would take to reach the ground is most nearly


1
Expert's answer
2021-02-23T10:02:45-0500
"t=\sqrt{\frac{2h}{g}}\\t=\sqrt{\frac{2(20)}{9.8}}=2.0\ s"